確率解析とその周辺

平成 16 年度科学研究費補助金基盤研究 (A)(1) 「確率論の総合的研究」 (研究代表者：重川一郎，課題番号： 14204008) による標記の研究集会を以下の要領で開催致しますのでご案内申し上げます．

場所：大阪大学基礎工学部シグマホールディスプレイ室

期日： 2005 年 1 月 7 日 (金) − 1 月 9 日 (日)

Program

January 7

13:20–14:10 Ichiro Shigekawa (Kyoto University)

Schrödinger operators on the Wiener space 14:20–15:10 Kazumasa Kuwada (Kyoto University)

Laplace approximation for stochastic line integrals 15:40–16:30 Minoru W. Yoshida (The University of Electro-Communications)

Homogenization of diffusions on the lattice Z^{d} with periodic drift coefficients; Application of logarithmic Sobolev inequality (with S. Albeverio, M. S. Bernabei and M. Röckner)

January 8

10:00–10:50 Shigeki Aida (Osaka University)

Semiclassical limit of the lowest eigenvalue of Schrödinger operators in path spaces over Riemannian manifolds

11:00–11:50 Zdzisław Brzeźniak (The University of Hull)

Stochastic nonlinear beam equation 13:20–14:10 Naoki Heya (The University of Tokyo)

Hypoellipticity in infinite dimensions 14:20–15:10 Kazuhiro Kuwae (Kumamoto University)

Kato class measures under heat kernel estimate (with M. Takahashi) 15:40–16:30 Short Communications

January 8

10:00–10:50 Masanori Hino (Kyoto University)

A trace theorem for Dirichlet forms on fractals (with T. Kumagai) 11:00–11:50 Zdzisław Brzeźniak (The University of Hull)

Asymptotic compactness and absorbing sets for stochastic Burgers equations driven by space time white noise and for some 2D stochastic Navier-Stokes equations on certain unbounded domains

世話人：会田茂樹 (大阪大学基礎工学研究科) 重川一郎 (京都大学理学研究科)